### v1

# Intersections of blocks of cyclotomic Hecke algebras

Abstract.
Trinh and Xue have proposed a startling conjecture on intersections of blocks of cyclotomic Hecke algebras occurring in modular representation theory of finite reductive groups. We prove this conjecture for all exceptional type groups apart from $E_{8}$. We also propose several generalisations, to Suzuki and Ree groups, to non-rational Coxeter groups and even more generally to spetsial complex reflection groups, and confirm these in various cases.
Key words and phrases:
blocks of Hecke algebras, Harish-Chandra theory, complex reflection groups, spetses
2010 Mathematics Subject Classification:
20C08, 20C33
Acknowledgment by the first author. Research supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the Basic Research Financing (Horizontal support for all Sciences), National Recovery and Resilience Plan (Greece 2.0), Project Number: 15659, Project Acronym: SYMATRAL
1. Introduction
Cyclotomic Hecke algebras were introduced in [5] to explain observations in the modular representation theory of finite reductive groups in non-defining characteristic. In this paper we study their specialisations at roots of unity. Our research is motivated by a surprising conjecture of Trinh and Xue [31, Conj. 4.1], which we prove for all exceptional groups apart from a few cases in type $E_{8}$. We also give a proof in all cases where relative Weyl groups are cyclic.
Let $G$ be a finite reductive group with Weyl group $W$ and ${\mathcal{H}}(W)$ the classical generic Iwahori–Hecke algebra of $W$. The conjecture proposes and the checked examples show that the $d$-cyclotomic Hecke algebras constructed in [5] from certain relative Weyl groups in $G$ “know” about the blocks of root-of-unity specialisations of ${\mathcal{H}}(W)$. We note that while the $d$-cyclotomic Hecke algebras were introduced to (conjecturally) explain the block distribution of $G$ at non-defining primes, the blocks of the cyclotomic Hecke algebras as well as those of ${\mathcal{H}}(W)$ are usually smaller than those of $G$, so this relation is mysterious even if one accepts that cyclotomic Hecke algebras do control the block theory of $G$.
Theorem 1.1.
The Conjecture 2.3 holds for all finite reductive groups of exceptional type, except possibly for type $E_{8}$ when $d\in\{3,4,6\}$.
For $E_{8}$ and $d=3,4,6$ we have obtained approximate block decompositions that are in accordance with Conjecture 2.3.
We also propose a natural generalisation of the conjecture in [31] to arbitrary finite Coxeter groups as well as to the setting of spetses and prove:
Theorem 1.2.
The generalised Conjecture 4.1 holds for the non-crystallographic Coxeter groups as well as for the spetsial primitive complex reflection groups $G_{4},G_{6},G_{8}$ and $G_{24}$.
We also provide partial results for $G_{33}$.
In Section 2 we introduce the setting, state the conjecture of Trinh and Xue, collect some tools and describe our general approach. We then verify the conjecture for groups of exceptional Lie type in Section 3, proving Theorem 1.1. In Section 4 we formulate an extension of the conjecture to the spetsial setting and prove Theorem 1.2.
2. Blocks in Harish-Chandra series
2.1. Cyclotomic Hecke algebras
Let ${\mathbf{G}}$ be a simple algebraic group over an algebraically closed field of characteristic $p$ with a Frobenius endomorphism $F:{\mathbf{G}}\to{\mathbf{G}}$ defining an ${\mathbb{F}}_{q}$-structure, and $G:={\mathbf{G}}^{F}$ the group of $F$-fixed points, a finite reductive group. While most of what is explained below holds more generally for ${\mathbf{G}}$ just being connected reductive, for simplicity we stick to simple groups. We write ${\operatorname{Uch}}(G)$ for the set of unipotent characters of $G$.
Let $e\geq 1$ be an integer such that the cyclotomic polynomial $\Phi_{e}$ divides the order polynomial of $({\mathbf{G}},F)$. Recall that the centralisers of $e$-tori of ${\mathbf{G}}$ are called $e$-split Levi subgroups. These are certain $F$-stable Levi subgroups (of not necessarily $F$-stable parabolic subgroups) of ${\mathbf{G}}$ (see e.g. [16, 3.5.1]). A unipotent character $\lambda\in{\operatorname{Uch}}(G)$ is called $e$-cuspidal if ${}^{*}\!R_{\mathbf{L}}^{\mathbf{G}}(\lambda)=0$ for all proper $e$-split Levi subgroups of ${\mathbf{G}}$. Here, ${}^{*}\!R_{\mathbf{L}}^{\mathbf{G}}$ denotes Lusztig restriction. A unipotent $e$-cuspidal pair of ${\mathbf{G}}$ is a pair $({\mathbf{L}},\lambda)$ consisting of an $e$-split Levi subgroup ${\mathbf{L}}$ of ${\mathbf{G}}$ and an $e$-cuspidal unipotent character $\lambda\in{\operatorname{Uch}}({\mathbf{L}}^{F})$. These play a fundamental role in the $\ell$-modular representation theory of $G$ for primes $\ell$ dividing $\Phi_{e}(q)$, the $e$th cyclotomic polynomial evaluated at $q$, in that they serve to parametrise the unipotent $\ell$-blocks of $G$. Since in this paper we will solely be interested in unipotent characters, we will just speak of $e$-cuspidal pairs when we mean unipotent $e$-cuspidal pairs.
Associated to any $e$-cuspidal pair $({\mathbf{L}},\lambda)$ is its relative Weyl group defined as $W_{\mathbf{G}}({\mathbf{L}},\lambda):=N_{G}({\mathbf{L}},\lambda)/{\mathbf{L}}^{F}$. If the ambient group ${\mathbf{G}}$ is clear from the context, we will also denote it by $W({\mathbf{L}},\lambda)$. It is known by inspection that $W({\mathbf{L}},\lambda)$ is always a finite complex reflection group over an abelian number field $K$, irreducible if ${\mathbf{G}}$ is simple (see [5, §3] and [6, Tab. 3]) Let ${\mathcal{H}}(W({\mathbf{L}},\lambda),{\mathbf{u}})$ be the generic Hecke algebra of $W({\mathbf{L}},\lambda)$ over ${\mathbb{Z}}[{\mathbf{u}}^{\pm 1}]$, for ${\mathbf{u}}$ a set of algebraically independent parameters (as defined for example in [8, 1.5]). In [5] and [21] we single out a particular 1-parameter specialisation ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ of ${\mathcal{H}}(W({\mathbf{L}},\lambda),{\mathbf{u}})$ with respect to a specialisation map ${\mathbb{Z}}[{\mathbf{u}}^{\pm 1}]\to K[x^{\pm\frac{1}{2}}]$ depending on ${\mathbf{G}}$ and $({\mathbf{L}},\lambda)$, the cyclotomic Hecke algebra associated to $({\mathbf{L}},\lambda)$ in ${\mathbf{G}}$. Conjecturally, the specialisation of ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ at $x\mapsto q$ should be the endomorphism algebra of the complex defining Lusztig induction $R_{\mathbf{L}}^{\mathbf{G}}(\lambda)$, see [5, ($d$-HV3)]; this has been shown in a number of situations by Lusztig, Digne, Michel, Rouquier and Dudas (see [31, Thm 3.6] for a list of known instances), but is open in general. Nevertheless, all numerical consequences implied by this have been shown to hold [24].
Now the Hecke algebra ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ (conjecturally) carries a canonical symmetrising form, which should, in particular, specialise to the natural form on the endomorphism algebra of $R_{\mathbf{L}}^{\mathbf{G}}(\lambda)$. Existence of this form is known when $W({\mathbf{L}},\lambda)$ is a Weyl group and has been established case-by-case for all but a handful of primitive complex reflection groups: for the imprimitive groups in [4, 25], and for
  • the group $G_{4}$ in [26, 27, 3] (3 independent proofs);
  • the groups $G_{5},G_{6},G_{7},G_{8}$ in [3];
  • the group $G_{12}$ in [26];
  • the group $G_{13}$ in [2];
[…]
  • the groups $G_{22},G_{24}$ in [26].
Let $\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ denote the set of characters of the irreducible representations of ${\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ over a splitting field. For $\chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ we denote its Schur elements with respect to this form by $S_{\chi}$, an element of $K[x^{\pm\frac{1}{m}}]$ for some suitable $m\geq 1$. These Schur elements have been computed, assuming the form exists, in all cases (see [24] and the references therein).
Let ${\operatorname{Uch}}(G,({\mathbf{L}},\lambda)):=\{\chi\in{\operatorname{Uch}}(G)\mid\langle R_{\mathbf{L}}^{\mathbf{G}}(\lambda),\chi\rangle\neq 0\}$ denote the $e$-Harish-Chandra series of $G$ above the $e$-cuspidal pair $({\mathbf{L}},\lambda)$. We then have:
Theorem 2.1.
For each $e$-cuspidal pair $({\mathbf{L}},\lambda)$ there is a bijection
$\Psi_{{\mathbf{L}},\lambda}:\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))\longrightarrow{\operatorname{Uch}}(G,({\mathbf{L}},\lambda))$
satisfying
()( ∗ ) $\Psi_{{\mathbf{L}},\lambda}(\chi)(1)=\pm\lambda(1)\,|G:L|_{p^{\prime}}/\,S_{\chi}(q)\qquad\text{for all $\ \chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$}.$
The first part of the above theorem is [6, Thm 3.2]. The second part was conjectured in [5, ($d$-HV6)] and the proof was completed in [24, Prop. 7.1], using earlier work of several authors on Schur elements (as explained in the paragraph before [24, Prop. 7.1]).
[…]
Remark 2.2.
For $e=1$ the bijections $\Psi_{{\mathbf{L}},\lambda}$ come from the identification of the cyclotomic Hecke algebras ${\mathcal{H}}(W({\mathbf{L}},\lambda),q)$ as endomorphism algebras of Harish-Chandra induced modules and provide the standard labelling of unipotent characters of $G$. For $e>1$ the bijections should be induced by the cited conjectural relation to endomorphism algebras as well; this has not been proved, and condition ()(*)( ∗ ) does not, in general, determine $\Psi_{{\mathbf{L}},\lambda}$ uniquely though.
2.2. The conjecture of Trinh and Xue
We can now describe the conjecture of Trinh and Xue. For this, consider a second integer $d\geq 1$ such that $\Phi_{d}$ also divides the order polynomial of $({\mathbf{G}},F)$, and let $({\mathbf{M}},\mu)$ be a $d$-cuspidal pair of ${\mathbf{G}}$, with associated cyclotomic Hecke algebra ${\mathcal{H}}(W({\mathbf{M}},\mu),x)$. Let $\zeta_{d},\zeta_{e}$ be, respectively, primitive $d$th and $e$th roots of unity, and consider the specialisations ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$ and ${\mathcal{H}}(W({\mathbf{M}},\mu),\zeta_{e})$. These will, in general, not be semisimple, so will induce some non-trivial block partitions on $\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ and $\operatorname{Irr}({\mathcal{H}}(W({\mathbf{M}},\mu),x))$ respectively. Note that when $e=1$, the blocks of ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$ are related to the $\ell$-block structure of $G$ for primes $\ell$ dividing $\Phi_{d}(q)$ (see e.g. [15, §4]), but no such relation is known (to us) when $e>1$.
Trinh and Xue want to relate these block partitions by comparing their images under $\Psi$ in the corresponding Harish-Chandra series of $G$. Note that these will depend on the choice of complete HC-parametrisation $\Psi$. Concretely, [31, Conj. 4.1] predicts the following (denoted properties (I) and (II) in loc. cit.):
Conjecture 2.3 (Trinh–Xue).
There exists a choice of complete HC-parametrisation $\Psi=\{\Psi_{{\mathbf{L}},\lambda}\}$ such that for any $e$-cuspidal pair $({\mathbf{L}},\lambda)$ and $d$-cuspidal pair $({\mathbf{M}},\mu)$ the intersection of Harish-Chandra series
$\Psi_{{\mathbf{L}},\lambda}(\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x)))\ \cap\ \Psi_{{\mathbf{M}},\mu}(\operatorname{Irr}({\mathcal{H}}(W({\mathbf{M}},\mu),x)))$
admits a partition $\bigsqcup B_{i}$ such that for all $i$, $\Psi_{{\mathbf{L}},\lambda}^{-1}(B_{i})$ is a block of ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$ and $\Psi_{{\mathbf{M}},\mu}^{-1}(B_{i})$ is a block of ${\mathcal{H}}(W({\mathbf{M}},\mu),\zeta_{e})$. In particular, this defines a bijection between those blocks of the two algebras appearing in this intersection.
This claim has been verified in [31, Thm 3] for groups of type $\operatorname{GL}_{n}$ when $d,e$ are coprime, and in [31, Cor. 5 and §8] certain numerical consequences are checked for exceptional type groups. Of course the conjecture makes sense in a more general setting where ${\mathbf{G}}$ is just assumed to be connected reductive, but it easily reduces to the case of simple ${\mathbf{G}}$ as unipotent characters and Lusztig induction behave well with respect to central products. As we will see below, the conjecture also makes sense and holds when $F$ is not a Frobenius but a Steinberg map, so $G$ is a Suzuki or Ree group.
2.3. Verifying the conjecture
We start out by making some general remarks. The following observation is valid for any ${\mathbf{G}}$:
Lemma 2.4.
Assume that $({\mathbf{M}},\mu)=({\mathbf{G}},\mu)$ is a $d$-cuspidal pair of ${\mathbf{G}}$. Then Conjecture 2.3 holds for any $e$-cuspidal pair $({\mathbf{L}},\lambda)$, independently of the chosen HC-parametrisation.
Proof.
If $({\mathbf{M}},\mu)=({\mathbf{G}},\mu)$, that is, if $\mu$ is a $d$-cuspidal unipotent character of $G$, then $W({\mathbf{G}},\mu)=1$, so $\operatorname{Irr}({\mathcal{H}}(W({\mathbf{M}},\mu),x)=\{1\}$ is a singleton and hence consists of a single block. It follows by [6, Prop. 2.4] that the degree polynomial of $\mu$ is divisible by the full $\Phi_{d}$-part of the order polynomial of $({\mathbf{G}},F)$. Let $({\mathbf{L}},\lambda)$ be an $e$-cuspidal pair of ${\mathbf{G}}$ with $\chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ such that $\Psi_{{\mathbf{L}},\lambda}(\chi)=\mu$. Then by Theorem 2.1,
$\mu(1)=\Psi_{{\mathbf{L}},\lambda}(\chi)(1)=\lambda(1)|G:L|_{p^{\prime}}/S_{\chi}(q)$
is divisible by the full $\Phi_{d}$-part of the order polynomial of $({\mathbf{G}},F)$. Now the degree polynomial of $\lambda$ divides the order polynomial of $({\mathbf{L}},F)$, so the Schur element $S_{\chi}$ must be prime to $\Phi_{d}$. That is, its specialisation $S_{\chi}(\zeta_{d})$ at $x=\zeta_{d}$ of $S_{\chi}$ does not vanish. But then by general properties of symmetric algebras [17, Thm 7.5.11], $\chi$ lies in a singleton block of ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$, so Conjecture 2.3 holds for the pair $({\mathbf{L}},\lambda),({\mathbf{M}},\mu)$. ∎
We also have a general result at the opposite extreme from cuspidal characters. We know that $\chi,\psi\in\operatorname{Irr}({\mathcal{H}})$ lie in the same block if and only if the values of their central characters $\omega_{\chi},\omega_{\psi}$ agree on all central elements of ${\mathcal{H}}$ [17, Lemma 7.5.10]. In particular, if $\chi,\psi\in\operatorname{Irr}({\mathcal{H}})$ lie in the same block, then the values of their central characters $\omega_{\chi},\omega_{\psi}$ agree on a generator of the centre of the braid group. This defines a (possibly coarser) partition of $\operatorname{Irr}({\mathcal{H}})$ that we call the coarse blocks.
Proposition 2.5.
Assume that $({\mathbf{L}},1)$, $({\mathbf{M}},1)$ are respectively $e$-cuspidal and $d$-cuspidal pairs of ${\mathbf{G}}$. Then (assuming [14, Conj. 9.1] holds) Conjecture 2.3 for $({\mathbf{L}},1)$, $({\mathbf{M}},1)$ is compatible with the coarse blocks.
Proof.
By [14, Cor. 9.5], assuming [14, Conj. 9.1], for any $\rho$ in the $e$-Harish-Chandra series of ${\mathbf{G}}^{F}$ above $({\mathbf{L}},1)$ the Frobenius $F^{\delta}$ acts by a scalar
$c_{{\mathbf{L}},1}\lambda_{\gamma}q^{-\delta(A_{\gamma}+a_{\gamma})/e}$
on the $\rho$-isotypic part of a suitable Deligne–Lusztig variety attached to $({\mathbf{L}},1)$, where $A_{\gamma}$ is the degree in $x$ of the degree polynomial of $\gamma$, $a_{\gamma}$ is the power of $x$ dividing it, and $\lambda_{\gamma}$ is the Frobenius eigenvalue of $\gamma$ and $c_{{\mathbf{L}},1}$ is a constant only depending on $({\mathbf{L}},1)$. Thus the corresponding central element of the cyclotomic Hecke algebra ${\mathcal{H}}(W({\mathbf{L}},1),x)$ acts by
$c_{{\mathbf{L}},1}\lambda_{\gamma}x^{-\delta(A_{\gamma}+a_{\gamma})/e}$
in the irreducible representation labelled by $\Psi_{{\mathbf{L}},1}(\gamma)$, and hence by the scalar
$c_{{\mathbf{L}},1}\lambda_{\gamma}\exp\big{(}-2\pi i\delta(A_{\gamma}+a_{\gamma})/(de)\big{)}$
in the specialisation ${\mathcal{H}}_{e}:={\mathcal{H}}(W({\mathbf{L}},1),\zeta_{d})$. Comparing with the corresponding scalars for the characters in the $d$-Harish-Chandra series above $({\mathbf{M}},1)$ we see that two characters of ${\mathcal{H}}_{e}$ lie in the same coarse block if and only if their images under $\Psi_{{\mathbf{M}},1}^{-1}\circ\Psi_{{\mathbf{L}},1}$ lie in the same coarse block for ${\mathcal{H}}_{d}:={\mathcal{H}}(W({\mathbf{M}},1),\zeta_{e})$. ∎
This applies in particular if $d$ is a regular number for $W$ (see e.g. [9, Def. 3.1]), and ${\mathbf{L}},{\mathbf{M}}$ are the centralisers of respectively Sylow $e$-tori and Sylow $d$-tori of ${\mathbf{G}}$.
2.4. The cyclic case
Assume $W({\mathbf{L}},\lambda)$ is cyclic, of order $m$. Then ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ is generated by a single element $T$ satisfying a relation $\prod_{i=1}^{m}(T-\gamma_{i}x^{a_{i}})=0$ for suitable roots of unity $\gamma_{i}$ and half-integral exponents $a_{i}$ (see [5, §2]). Thus, all irreducible characters of ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ (over a splitting field) are linear, and there are exactly $m$ of them, defined by
()( † ) $\chi_{i}:T\mapsto\gamma_{i}x^{a_{i}},\qquad i=1,\ldots,m.$
In the specialisation ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$, two characters $\chi_{i},\chi_{j}$ lie in the same block if and only if their specialisations agree, so if and only if $\gamma_{i}\zeta_{d}^{a_{i}}=\gamma_{j}\zeta_{d}^{a_{j}}$. Hence in this case the determination of blocks is immediate from the parameters defining the cyclotomic Hecke algebra.
Corollary 2.6.
Assume $d,e$ are both regular. Let $({\mathbf{L}},1)$, $({\mathbf{M}},1)$ be respectively $e$-cuspidal and $d$-cuspidal pairs such that both relative Weyl groups $W({\mathbf{L}},1)$ and $W({\mathbf{M}},1)$ are cyclic. Then (assuming [5, ($d$-HV3)] holds) Conjecture 2.3 holds for the intersection of the corresponding two Harish-Chandra series.
Proof.
By the remark after [9, Lemma 5.11] in this case ${\mathcal{H}}$ is generated by the element $F$ and hence by what we observed just above, coarse blocks agree with blocks, so the claim follows by Proposition 2.5. ∎
A general formula for values of characters of cyclotomic Hecke algebras on central elements is given in [7, Prop. 6.15].
2.5. Ennola duality
There is a further simplification coming from Ennola duality (see [6, (1.5), (1.21), Thm 3.2]). For integers $e\geq 1$ let
$e^{\prime}:=\begin{cases}2e&\text{if $e\equiv 1\pmod{2}$},\\ e/2&\text{if $e\equiv 2\pmod{4}$},\\ e&\text{if $e\equiv 0\pmod{4}$}.\end{cases}$
Then we have $\Phi_{e}(-x)=\pm\Phi_{e^{\prime}}(x)$ (where the minus sign occurs only when $e\in\{1,2\}$). Now associated to $({\mathbf{G}},F)$ is its Ennola dual group $({\mathbf{G}},F^{\prime})$ whose complete root datum is obtained by formally replacing $F$ by $-F$. We have the following:
Lemma 2.7.
Let ${\mathbf{G}}$ be as above and $d,e\geq 1$.
  1. (a)
    If Conjecture 2.3 holds for $({\mathbf{G}},F)$ at $(d,e)$ it also holds for $({\mathbf{G}},F^{\prime})$ at $(d^{\prime},e^{\prime})$.
  2. (b)
    If $W$ contains $-\operatorname{Id}$ then Conjecture 2.3 holds for $(e,e^{\prime})$ for all $e\geq 1$ with respect to a suitable complete HC-parametrisation.
Proof.
The unipotent characters of the finite reductive group ${\mathbf{G}}^{F^{\prime}}$ are in bijection with those of ${\mathbf{G}}^{F}$, where the degree functions are changed by replacing $x$ by $-x$ [6, Thm 3.3]. Further, for any $F$-stable maximal torus ${\mathbf{T}}$ of ${\mathbf{G}}$ there exists an $F^{\prime}$-stable maximal torus ${\mathbf{T}}^{\prime}$ of ${\mathbf{G}}$ whose order polynomial is obtained from the one of ${\mathbf{T}}$ by replacing $x$ by $-x$. Also, the relative Weyl groups of ${\mathbf{T}}$ and ${\mathbf{T}}^{\prime}$ agree and thus so do their generic Hecke algebras. More generally, the $e$-cuspidal pairs of $({\mathbf{G}},F)$ are in bijection with the $e^{\prime}$-cuspidal pairs of $({\mathbf{G}},F^{\prime})$, with equal relative Weyl groups and thus Hecke algebras. It is part of the conjectures in [5] that the cyclotomic specialisations for cuspidal pairs of $({\mathbf{G}},F^{\prime})$ are obtained from the ones for $({\mathbf{G}},F)$ by replacing $x$ by $-x$. Then clearly the $\zeta_{d}$-blocks of the former are mapped to $\zeta_{d^{\prime}}$-blocks of the latter. So if Conjecture 2.3 holds for $({\mathbf{G}},F)$ and $(d,e)$, it will automatically hold for $({\mathbf{G}},F^{\prime})$ and $(d^{\prime},e^{\prime})$.
If $W$ contains $-\operatorname{Id}$ then $({\mathbf{G}},F)$ and $({\mathbf{G}},F^{\prime})$ are isomorphic and hence their sets of unipotent characters, cuspidal pairs and relative Weyl groups may be identified. Let $({\mathbf{L}},\lambda)$ be an $e$-cuspidal pair of $({\mathbf{G}},F)$ with cyclotomic Hecke algebra ${\mathcal{H}}:={\mathcal{H}}(W({\mathbf{L}},\lambda),x)$. Then clearly, the specialisation of ${\mathcal{H}}$ at $x$ a primitive $e^{\prime}$th root of unity is identical to the specialisation of ${\mathcal{H}}$ at $-x$ a suitable primitive $e$th root of unity, by the definition of $e^{\prime}$. Hence, the blocks of both are the same and choosing a suitable HC-parametrisation the claim follows. ∎
2.6. Algorithmic determination of blocks
We now describe how we determined blocks of specialised Hecke algebras. Let $W$ be a complex reflection group and let ${\mathcal{H}}(W)$ denote a cyclotomic Hecke algebra over $R:={\mathbb{C}}[x,x^{-1}]$, where $y^{z}=x$ with $z=|Z(W)|$. It is known that $\operatorname{Frac}(R)$ is a splitting field for ${\mathcal{H}}(W)$. A representation $(V,\rho)$ of ${\mathcal{H}}(W)$ is a ${\mathcal{H}}(W)$-module $V$ which is finite-dimensional free as $R$-module, with structural morphism $\rho:{\mathcal{H}}(W)\to\operatorname{End}(V)$. Let $\zeta_{e}\in{\mathbb{C}}$ be a primitive $e$th root of unity and let $\zeta\in{\mathbb{C}}$ with $\zeta^{z}=\zeta_{e}$. We consider the ring homomorphism $\theta:R\to{\mathbb{C}}$ defined by $\theta(y)=\zeta$, so $\theta(x)=\zeta_{e}$, and let ${\mathcal{H}}_{e}(W)$ denote the corresponding specialised Hecke algebra over ${\mathbb{C}}$ (note that this depends on the choice of $\zeta$). If $\chi\in\operatorname{Irr}(W)$, then we denote by $S_{\chi}$ its Schur element in ${\mathcal{H}}(W)$ and we call defect of $\chi$ with respect to $\theta$ the multiplicity of $\zeta$ as a root of $S_{\chi}$.
If $W$ is a cyclic group, then the blocks of ${\mathcal{H}}_{e}(W)$ are easy to compute, because two characters $\chi,\psi$ lie in the same block if and only if $\theta(\chi(h))=\theta(\psi(h))$ for all $h\in{\mathcal{H}}(W)$, see Section 2.4 above.
The blocks of ${\mathcal{H}}_{e}(W)$ for the exceptional finite Coxeter groups $W$ can be found in [17, App. F], while for many cases of groups of rank 2 they have been calculated in [12]. The algorithm used in the latter reference is similar to the one we applied to the rest of the cases. This algorithm is heuristic and it demands that we apply all of the following arguments to determine mostly when two characters do not lie in the same block. In most instances, we explicitly computed the decomposition matrix $D=(d_{\chi,\varphi})_{\chi\in\operatorname{Irr}(W),\varphi\in\operatorname{Irr}({\mathcal{H}}_{e}(W))}$. For $G_{25}$ a couple of cases were taken from [30].
  1. (1)
    A character $\chi$ is alone in its block if and only if $\theta(S_{\chi})\neq 0$ ([17, Thm 7.5.11] and [18, Prop. 4.4]).
  2. (2)
    Any representation of dimension $1$ is irreducible. A representation of dimension $2$ is irreducible, unless it has a 1-dimensional subrepresentation. A representation of dimension $3$ is irreducible unless it or its transpose has a 1-dimensional subrepresentation. Dimension $1$ subreprepresentations correspond to common eigenvectors of the matrices representing the generators of ${\mathcal{H}}(W)$, so they are easy to find.
  3. (3)
    In general, assume that $(V,\rho)$, $(V^{\prime},\rho^{\prime})$ are simple modules of ${\mathcal{H}}(W)$. Then $(V,\theta\circ\rho)$, $(V^{\prime},\theta\circ\rho^{\prime})$ are modules of the specialised algebra ${\mathcal{H}}_{e}(W)$ (where by $\theta\circ\rho$ we mean that we apply $\theta$ to the entries of the matrix $\rho(h)$ for all $h\in{\mathcal{H}}(W)$). If $(V^{\prime},\theta\circ\rho^{\prime})$ is isomorphic to a submodule of $(V,\theta\circ\rho)$, then there exists a (rectangular) matrix $M$ such that $(\theta\circ\rho^{\prime}(s))\cdot M=M\cdot(\theta\circ\rho(s))$ for every generator $s$ of ${\mathcal{H}}(W)$ [30]. We only used this argument for the groups $G_{25}$ and $G_{26}$.
  4. (4)
    We have that $\chi,\psi$ are in the same block if and only if $\theta(\omega_{\chi}(z))=\theta(\omega_{\psi}(z))$ for all $z\in Z({\mathcal{H}}(W))$, where $\omega_{\chi}$ denotes the central character of $\chi$ [17, Lemma 7.5.10]. A known element of $Z({\mathcal{H}}(W))$ is the generator $\boldsymbol{\beta}$ of the centre of the corresponding braid group (proved to be cyclic by Bessis [1, Thm 12.8]). So if $\theta(\omega_{\chi}(\boldsymbol{\beta}))\neq\theta(\omega_{\psi}(\boldsymbol{\beta}))$, then $\chi,\psi$ are not in the same block.
  5. (5)
    Let $\varphi\in\operatorname{Irr}({\mathcal{H}}_{e}(W))$ and let $P(x)$ be a polynomial that is divisible by $S_{\chi}$ for all $\chi$ such that $d_{\chi,\varphi}\neq 0$. Then (see [12, Lemma 6])
    $\sum_{\chi\in\operatorname{Irr}(W)}d_{\chi,\varphi}\cdot\theta({P(x)}/{S_{\chi}})=0.$
    The fraction ${P(x)}/{S_{\chi}}$ could simply be the generic degree $D_{\chi}$, but when a character is not of maximal defect in ${\mathcal{H}}(W)$ (that is, $\theta(D_{\chi})=0$), it is convenient to use the least common multiple of the Schur elements of the characters in it.
Using the above arguments we were able to explicitly compute the blocks in all cases except for the groups $G_{31},G_{32}$ (which occur for $d=4$, respectively $d=3,6$ in $E_{8}$) and for $G_{33}$. In these three cases, using the above algorithm, we were able to split the characters into “approximate blocks”, but these could still be unions of blocks (our algorithm determining when characters are not in the same block, but not when they are). Note that there is also data missing regarding the representations of the corresponding Hecke algebras, which made our job more difficult, especially for $G_{31}$.
Remark 2.8.
After numerous calculations of blocks for cyclotomic Hecke algebras, it is worth pointing out the following facts:
  1. (1)
    We verified that the “defect conjecture” [13, Conj. 2] holds in every case, that is, all characters in a block have the same defect.
  2. (2)
    The irreducible representations of ${\mathcal{H}}_{e}(W)$ are liftable for all rank $2$ groups (this is what is called the existence of an optimal basic set in [12]), but not for rank $3$ or higher.
  3. (3)
    In most of the cases, the converse of the last statement of (4) above holds, that is, if $\theta(\omega_{\chi}(\boldsymbol{\beta}))=\theta(\omega_{\psi}(\boldsymbol{\beta}))$, then $\chi,\psi$ are in the same block. However, we have found a few exceptions (even in rank $2$ groups), so this does not hold in general. (Examples are given by ${\mathcal{H}}(G_{5})$ with parameters $1,x,x^{2}$ for $x=\zeta_{4}$, and ${\mathcal{H}}(G_{22})$ with parameters $1,x^{2}$ for $x=\zeta_{5}$.)
All blocks we computed can be found in the github repository [32].
3. Blocks in the exceptional types
We now discuss Conjecture 2.3 for the various types of exceptional groups, by increasing rank. Our notation for unipotent characters follows [10] which is also the one used in the computer algebra package Chevie[28] from which all data are taken.
Note that we do not (need to) consider groups of type ${}^{2}\!E_{6}$, since these are obtained from the untwisted groups $E_{6}$ by Ennola duality (see Lemma 2.7 above). Since Conjecture 2.3 is symmetric in $d$ and $e$ and moreover vacuously satisfied when $d=e$, we may and will assume from now on that $d<e$, and also that both ${\mathbf{L}},{\mathbf{M}}$ are proper Levi subgroups of ${\mathbf{G}}$ (by Lemma 2.4).
It is our understanding that in [31, §8] it is shown for types $G_{2}$ and $F_{4}$ that the block sizes of the specialised cyclotomic algebras do match up as predicted by Conjecture 2.3, but without making explicit a choice of HC-parametrisation and (hence) also without checking that the images of the blocks under $\Psi_{{\mathbf{L}},\lambda},\Psi_{{\mathbf{M}},\mu}$ do actually agree. We check this latter property using the data and the functionality of the Chevie-system [28].
3.1. Type $G_{2}$
We start with type $G_{2}$. The block sizes for the relevant algebras have already been determined in [31, 8.6]. Here we also consider the intersection of the corresponding sets of unipotent characters. There are two pairs of characters of $G_{2}(q)$ not distinguished by their degrees, namely the principal series characters $\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime}$ and the cuspidal characters $G_{2}[\theta],G_{2}[\theta^{2}]$ of $G_{2}(q)$, where as customary $\theta$ denotes a primitive third root of unity. For these, there exist several possible Harish-Chandra parametrisations. Now it turns out that $\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime}$ in all cases either lie in the same block or they both occur in a singleton, so for the verification of Conjecture 2.3 no ambiguity arises. On the other hand, for $G_{2}[\theta],G_{2}[\theta^{2}]$ a parametrisation via endomorpism algebras is fixed by [19, (7.3)] for $d=6$. Hence the only relevant choice is for $d=3$. We present a compatible HC-parametrisation in Table 1. Here the characters of ${\mathcal{H}}(W(G_{2}),x)$ are identified through their standard Chevie-names, the (linear) character for cyclotomic Hecke algebras for cyclic groups by the corresponding parameter as in ()($\dagger$)( † ). With this, we have:
Proposition 3.1.
Conjecture 2.3 holds for type $G_{2}$ with respect to some complete HC-parametrisation. The block intersections are displayed in Table 1 (up to Ennola duality).
Table 1. Block intersections for $G_{2}$
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6}\}\\ (1,3)&\{\phi_{1,0},\phi_{1,6},\phi_{2,2}\}\\ (1,6)&\{\phi_{1,0},\phi_{1,6},\phi_{2,1}\}\\ (3,6)&\{\phi_{1,0},G_{2}[\theta]\},\{\phi_{1,6},G_{2}[\theta^{2}]\}\\ \end{array}$
Here, we have omitted all singleton blocks.
Proof.
The relative Weyl groups are cyclic or of type $G_{2}$. The parameters for the cyclotomic Hecke algebra are known by the work of Lusztig or predicted in [5, Bem. 2.7] (see also the table compiled in [31, 8.6]). The decomposition matrices for specialisations of ${\mathcal{H}}(W(G_{2}),x)$ are given in [15, Tab. 7.3], the ones for the cyclic case were described in Section 2.3. The claim follows by direct computation. ∎
3.2. Type ${}^{3}\!D_{4}$
All unipotent characters of the triality groups ${}^{3}\!D_{4}(q)$ are uniquely determined by their degree polynomial and hence ()(*)( ∗ ) determines a unique complete HC-parametrisation.
Proposition 3.2.
Conjecture 2.3 holds for type ${}^{3}\!D_{4}$; the block intersections are displayed in Table 2 (up to Ennola duality).
Table 2. Block intersections for ${}^{3}\!D_{4}$
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6}\}\\ (1,3)&\{\phi_{1,0},\phi_{1,6},\phi_{2,2}\},\{\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{2,1}\}\\ (1,6)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6},\phi_{2,2}\}\\ (1,12)&\{\phi_{1,0},\phi_{1,6},\phi_{2,1}\}\\ (3,6)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6},\phi_{2,2},{}^{3}\!D_{4}[1]\}\\ (3,12)&\{\phi_{1,0},\phi_{1,6},\phi_{2,1}\}\\ \end{array}$
Here, again, we have omitted all singleton blocks.
Proof.
The relative Weyl groups are described in [6, Tab. 3], the parameters for the corresponding cyclotomic Hecke algebras are predicted in [5, Folg. 5.6] (see also the table compiled in [31, 8.8.1]). The blocks for non-cyclic $W$ are again to be found in [15, Tab. 7.4] for type $W(G_{2})$, and have been worked out in [12, 3.3.1] for the primitive complex reflection group $G_{4}$, respectively. ∎
3.3. Type $F_{4}$
We now turn to type $F_{4}$. Again, the block sizes for the relevant algebras have already been determined by Trinh and Xue [31, 8.7]. Here, we also compare the actual intersections of Harish-Chandra series. This group is particularly interesting as there are ten pairs of unipotent characters of $F_{4}(q)$ having equal degree, so for all of those are there potential choices in the HC-parametrisation.
Proposition 3.3.
Conjecture 2.3 holds for type $F_{4}$ with respect to some complete HC-parametrisation. The block intersections are displayed in Table 3.
Table 3. Block intersections for $F_{4}$
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{\phi_{1,0},\phi_{1,12}^{*},\phi_{1,24},\phi_{2,4}^{*},\phi_{2,16}^{*},\phi_{9,2},\phi_{9,6}^{*},\phi_{9,10},\phi_{6,6}^{*},\phi_{8,3}^{*},\phi_{8,9}^{*}\},\\ &\{\phi_{4,1},\phi_{4,7}^{*},\phi_{4,13}\},\{B_{2}[.1^{2}],B_{2}[.2],B_{2}[1^{2}.],B_{2}[2.]\}\\ (1,3)&\{\phi_{1,0},\phi_{1,12}^{*},\phi_{1,24},\phi_{2,4}^{*},\phi_{2,16}^{*},\phi_{4,8}\},\{\phi_{4,1},\phi_{4,7}^{*},\phi_{4,13},\phi_{8,3}^{*},\phi_{8,9}^{*},\phi_{16,5}\}\\ (1,4)&\{\phi_{1,0},\phi_{1,24},\phi_{4,8},\phi_{9,2},\phi_{9,10},\phi_{6,6}^{\prime},\phi_{12,4},\phi_{4,1},\phi_{4,13}\},\\ &\{\phi_{2,4}^{\prime},\phi_{2,16}^{\prime},\phi_{4,7}^{\prime}\},\{\phi_{2,4}^{\prime\prime},\phi_{2,16}^{\prime\prime},\phi_{4,7}^{\prime\prime}\},\{B_{2}[.1^{2}],B_{2}[1.1],B_{2}[2.]\}\\ (1,6)&\{\phi_{1,0},\phi_{1,24},\phi_{2,4}^{*},\phi_{2,16}^{*},\phi_{9,6}^{*},\phi_{12,4},\phi_{8,3}^{*},\phi_{8,9}^{*}\},\\ &\{B_{2}[.1^{2}],B_{2}[.2],B_{2}[1^{2}.],B_{2}[2.]\}\\ (1,12)&\{\phi_{1,0},\phi_{1,24},\phi_{6,6}^{\prime\prime},\phi_{4,1},\phi_{4,13}\},\{B_{2}[.1^{2}],B_{2}[1.1],B_{2}[2.]\}\\ (3,4)&\{\phi_{1,0},\phi_{1,24},\phi_{4,8},\phi_{4,1},\phi_{4,13},F_{4}^{2}[1]\},\{\phi_{2,4}^{\prime},\phi_{2,16}^{\prime},\phi_{4,7}^{\prime}\},\{\phi_{2,4}^{\prime\prime},\phi_{2,16}^{\prime\prime}\phi_{4,7}^{\prime\prime}\},\\ (3,6)&\{\phi_{1,0},\phi_{2,16}^{*},\phi_{8,3}^{*},F_{4}[\theta^{2}]\},\{\phi_{1,24},\phi_{2,4}^{*},\phi_{8,9}^{*},F_{4}[\theta]\},\\ (3,12)&\{\phi_{1,0},\phi_{4,13},F_{4}[\theta]\},\{\phi_{1,24},\phi_{4,1},F_{4}[\theta^{2}]\}\\ (4,8)&\{\phi_{1,0},\phi_{9,10},F_{4}[i]\},\{\phi_{1,24},\phi_{9,2},F_{4}[-i]\}\\ (4,12)&\{\phi_{1,0},\phi_{1,24},B_{2}[1.1]\},\{\phi_{4,1},B_{2}[.1^{2}],F_{4}[-i]\},\{\phi_{4,13},B_{2}[2.],F_{4}[i]\}\\ (8,12)&\{\phi_{1,0},F_{4}[-i]\},\{\phi_{1,24},F_{4}[i]\}\\ \end{array}$
Here, an entry $\phi_{d,b}^{*}$ stands for the two characters $\phi_{d,b}^{\prime}$ and $\phi_{d,b}^{\prime\prime}$. Again, we have omitted all singletons, as well as the pairs $(1,8),(3,8)$ where there is just one non-singleton block.
Proof.
The relative Weyl groups are described in [6, Tab. 1 and 3], the parameters for the corresponding Hecke algebras are predicted in [5, Tab. 8.1]. The relevant block decompositions are taken from [17, Tab. F.3] for $W(F_{4})$ and from [12, 3.3.2, 3.3.3] for the primitive complex reflection groups $G_{5}$ and $G_{8}$ occurring as relative Weyl groups for $d=3,6$ and $d=4$, respectively. ∎
3.4. Type $E_{6}$
For $G$ of type $E_{6}$, again most block sizes for the relevant algebras have been determined in [31, 8.7]. Here, we also compare the blocks to the intersections of Harish-Chandra series. The only unipotent characters not distinguished by their degree are the two cuspidal ones, so only for those is there a choice in the HC-parametrisation.
Proposition 3.4.
Conjecture 2.3 holds for type $E_{6}$ with respect to some complete HC-parametrisation.
In this case and from now on, we do not display the block intersections for reasons of space, but they can all be found in the github repository [32].
Proof.
The relative Weyl groups are described in [6, Tab. 1 and 3], the parameters for the corresponding Hecke algebras are predicted in [5, Tab. 8.1]. The relevant decomposition matrices are taken from [17, Tab. F.4] for $W(E_{6})$ and [12, §3.3] for the 2-dimensional groups. The decomposition matrices for $G_{25}$ were computed using the methods expounded in Section 2.6 (with just a couple of cases taken from [30]). ∎
3.5. Type $E_{7}$
The groups of type $E_{7}$ possess four pairs of unipotent characters not distinguished by their degrees, so only for those is there a choice in the HC-parametrisation.
Proposition 3.5.
Conjecture 2.3 holds for type $E_{7}$ with respect to a suitable complete HC-parametrisation.
Proof.
The relative Weyl groups are described in [6, Tab. 1 and 3], the parameters for the corresponding Hecke algebras are predicted in [5, Tab. 8.1]. The relevant blocks are taken from [17, Tab. F.5] for $W(E_{7})$ and [12, §3.3] for 2-dimensional groups. The decomposition matrices for $G_{26}$ were computed using the methods explained in Section 2.6. ∎
Again, we do not print the actual intersections. In Table 4 we collect the number of non-singleton blocks appearing in the intersection of Harish-Chandra series for the relevant pairs $(d,e)$.
Table 4. Numbers of non-trivial block intersections for $E_{7}$
$\begin{array}[]{c|ccccccccccccc}e&d=2&3&4&5&6&7&8&9&10&12&14&18\\ \hline\cr 1&4&8&4&6&6&2&8&2&5&4&2&3\\ 3&&&12&6&13&4&10&4&9&6&4&4\\ 4&&&&8&12&4&16&4&8&6&4&4\\ 5&&&&&9&4&8&4&8&4&3&3\\ 7&&&&&&&4&2&3&2&2&1\\ 8&&&&&&&&4&8&6&4&4\\ 9&&&&&&&&&3&4&1&4\\ 12&&&&&&&&&&&2&4\\ \end{array}$
3.6. Type $E_{8}$
For type $E_{8}$ we have not been able to determine the block decompositions completely for all relevant cyclotomic Hecke algebras. More precisely, the dimensions of the irreducible representations of the 4-dimensional reflection group $G_{32}$ (and thus of its Hecke algebra), occurring as relative Weyl group for the principal series when $d=3$ or $d=6$, are too large. The situation for the 4-dimensional group $G_{31}$, occurring when $d=4$, is even worse, as the data available in Chevie are rather incomplete. In these two situations we only obtained approximations to the actual block decompositions, see the remarks at the end of Section 2.6. These approximations turned out to be in accordance with the Trinh–Xue conjecture as well.
Proposition 3.6.
Conjecture 2.3 holds for type $E_{8}$, except possibly when one of the two Harish-Chandra series is a principal $d$-series with $d\in\{3,4,6\}$.
Proof.
The relative Weyl groups are described in [6, Tab. 1 and 3], the parameters for the corresponding Hecke algebras are predicted in [5, Tab. 8.1]. The relevant blocks are taken from [17, Tab. F.6] for $W(E_{8})$ and [12, §3.3] for 2-dimensional groups. ∎
3.7. Suzuki and Ree groups
We now consider a closely related setting. In this section let ${\mathbf{G}}$ be simple of type $B_{2},G_{2}$ or $F_{4}$ in characteristic $p=2,3,2$ respectively, and assume $F:{\mathbf{G}}\to{\mathbf{G}}$ is a Steinberg endomorphism which is not a Frobenius, so $G:={\mathbf{G}}^{F}$ is a Suzuki group ${}^{2}\!B_{2}(q^{2})$ or a Ree group ${}^{2}G_{2}(q^{2})$ or ${}^{2}\!F_{4}(q^{2})$, for $q^{2}$ an odd power of $p$. Here, one can still define $\Phi$-tori, which satisfy a Sylow theory, where now $\Phi$ is a cyclotomic polynomial over ${\mathbb{Q}}(\sqrt{p})$, see e.g. [16, 3.5.3]. Furthermore, the relative Weyl groups are again complex reflection groups, and there is a $\Phi$-Harish-Chandra theory for unipotent characters as in the case of Frobenius maps [6, §3A]. It is therefore not surprising that the following extension of Conjecture 2.3 holds:
Proposition 3.7.
The Suzuki and Ree groups satisfy the analogue of Conjecture 2.3 for all pairs of cyclotomic polynomials over ${\mathbb{Q}}(\sqrt{p})$.
Proof.
The relative Weyl groups $W_{\Phi}$ of Sylow $\Phi$-tori in this case are cyclic unless $G={}^{2}\!F_{4}(q^{2})$ and the roots of $\Phi$ have order dividing 8, see e.g. [16, Tab. 3.2]. The parameters of the corresponding cyclotomic Hecke algebras have been given in [19] for the Coxeter case and in [5, Tab. 8.1] otherwise. The decomposition numbers for the non-cyclic relative Weyl groups $G_{8}$ and $G_{12}$ are taken from [12]. ∎
4. Extension to the spetsial case
We consider an extension of Conjecture 2.3 to spetsial complex reflection groups. These include, in particular, all finite Coxeter groups. We refer to [8] for notions and general facts about spetses.
4.1. A generalised conjecture
Let $W$ be a spetsial (finite) complex reflection group on $V={\mathbb{C}}^{n}$ (see [23, §3]). Let moreover $\varphi\in\operatorname{GL}(V)$ be normalising $W$. Then attached to the spets ${\mathbb{G}}:=(V,W\varphi)$ there is a set of unipotent characters ${\operatorname{Uch}}({\mathbb{G}})$ together with a degree function
${\operatorname{Uch}}({\mathbb{G}})\longrightarrow{\mathbb{C}}[x],\quad\rho\mapsto\rho(1),$
and an action of $N_{\operatorname{GL}(V)}(W\varphi)$ on ${\operatorname{Uch}}({\mathbb{G}})$ (see [20] for the non-rational Coxeter groups, [22] and [24, Thm 8.4] for the non-real cases, and also [8] for explicit tables for primitive $W$). If $W$ is a Weyl group, ${\operatorname{Uch}}({\mathbb{G}})$ are just the (labels of the) unipotent characters of an associated finite reductive group of rational type $W\varphi$ together with their degree polynomials. For any root of unity $\zeta$ and $\Phi:=x-\zeta$, there is defined a subset of $\Phi$-cuspidal elements of ${\operatorname{Uch}}({\mathbb{G}})$. A (unipotent) $\Phi$-cuspidal pair of ${\mathbb{G}}$ is a pair $({\mathbb{L}},\lambda)$ where ${\mathbb{L}}$ is a $\Phi$-split Levi of ${\mathbb{G}}$ and $\lambda\in{\operatorname{Uch}}({\mathbb{L}})$ is $\Phi$-cuspidal. For $({\mathbb{L}},\lambda)$ a $\Phi$-cuspidal pair let $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ be the stabiliser of $\lambda$ in $W_{\mathbb{G}}({\mathbb{L}})$ (see [8, Ax. 4.3]). Attached to $({\mathbb{L}},\lambda)$ is also a Hecke algebra ${\mathcal{H}}(W_{\mathbb{G}}({\mathbb{L}},\lambda),x)$ in one parameter $x$ (a specialisation of the generic Hecke algebra of $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ for parameters specified in [22, 21, 8]) naturally isomorphic (via the specialisation $x\mapsto\zeta$ and Tits’ deformation theorem) to the group algebra of $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ over a splitting field. This Hecke algebra should carry a canonical symmetrising form with associated Schur elements denoted $S_{\chi}$, see Section 2.1. According to [8, Ax. 4.31], there is a partition
${\operatorname{Uch}}({\mathbb{G}})=\displaystyle\coprod{\operatorname{Uch}}({\mathbb{G}},({\mathbb{L}},\lambda))$
into $\Phi$-Harish-Chandra series, where the union runs through representatives of $W$-orbits of $\Phi$-cuspidal pairs $({\mathbb{L}},\lambda)$ for ${\mathbb{G}}$, and as in the case of finite reductive groups (Theorem 2.1), for any such $({\mathbb{L}},\lambda)$ there is a bijection
$\Psi_{{\mathbb{L}},\lambda}:\operatorname{Irr}({\mathcal{H}}(W_{\mathbb{G}}({\mathbb{L}},\lambda),x))\rightarrow{\operatorname{Uch}}({\mathbb{G}},({\mathbb{L}},\lambda)),\qquad\varphi\mapsto\rho_{\varphi},$
such that the exact analogue of ()(*)( ∗ ) holds. It thus makes sense to propose the following, where $({\mathbb{M}},\mu)$ is a $\Phi^{\prime}=x-\zeta^{\prime}$-cuspidal pair of ${\mathbb{G}}$:
Conjecture 4.1.
In the above setting, the intersection of $\Phi$- resp. $\Phi^{\prime}$-Hardish-Chandra series
$\Psi_{{\mathbb{L}},\lambda}(\operatorname{Irr}({\mathcal{H}}(W({\mathbb{L}},\lambda),x)))\ \cap\ \Psi_{{\mathbb{M}},\mu}(\operatorname{Irr}({\mathcal{H}}(W({\mathbb{M}},\mu),x)))$
admits a partition $\bigsqcup B_{i}$ such that $\Psi_{{\mathbb{L}},\lambda}^{-1}(B_{i})$ is a block of ${\mathcal{H}}(W({\mathbb{L}},\lambda),\zeta^{\prime})$ and $\Psi_{{\mathbb{M}},\mu}^{-1}(B_{i})$ is a block of ${\mathcal{H}}(W({\mathbb{M}},\mu),\zeta)$ for all $i$, thus defining a bijection between those blocks of the two algebras appearing in this intersection.
Note that for $W$ a Weyl group this specialises to Conjecture 2.3.
It seems reasonable to expect that arguments similar to those used in [31, §6] might be applicable to show this for the infinite series of imprimitive spetsial groups.
Remark 4.2.
  1. (a)
    Observe that the case of Conjecture 4.1 where both relative Weyl groups are cyclic again holds by the analogue of Corollary 2.6.
  2. (b)
    There is again an Ennola phenomenon: If the centre of $W$ (assumed to be irreducible) is cyclic of order $z$, then replacing $x$ by $\exp(2\pi i/z)x$ induces a permutation on the set of unipotent characters as well as on the set of associated cyclotomic Hecke algebras (see [8, 4.2.6]) and hence of their block distributions, meaning that it suffices to check Conjecture 4.1 for just one representative in each orbit under this Ennola permutation (generalising Lemma 2.7(a)).
In the sequel we verify Conjecture 4.1 for some exceptional cases.
4.2. The non-crystallographic Coxeter groups
The first interesting case for Conjecture 4.1 is of course that of non-crystallographic Coxeter groups, that is, of types $I_{2}(m)$ with $m\geq 3$, $H_{3}$ and $H_{4}$.
Proposition 4.3.
Conjecture 4.1 holds for type $I_{2}(m)$, $m\geq 3$.
Proof.
The proper Levi subgroups of $I_{2}(m)$ are of maximal tori or of type $A_{1}$, thus the only proper $d$-cuspidal pairs are of type $({\mathbb{T}},1)$, for ${\mathbb{T}}$ a maximal torus containing a Sylow $d$-torus. The Sylow $d$-tori and hence their relative Weyl groups are cyclic unless $d\in\{1,2\}$, where $d=2$ only occurs when $m$ is even. Thus, by Corollary 2.6 for checking the conjecture we may assume that $d\leq 2$, and then by Ennola duality that $d=1$.
First assume $m$ is odd and $d=1$. We need to consider $\Phi=x-\zeta$-series for $\zeta$ some $m$th root of unity. Let $\zeta_{m}=\exp(2\pi i/m)$. By the action of Galois automorphisms we may assume $\zeta=\zeta_{m}^{m/e}$ for some divisor $e>2$ of $m$, a primitive $e$th root of unity. According to [15, Tab. 7.2] the Iwahori–Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,-1))$ specialised at $x=\zeta$ has just one non-singleton block, containing the trivial and the sign representation plus the representation denoted $\boldsymbol{\sigma}_{e}$. On the other hand, the Hecke algebra for the principal $\Phi$-series is ${\mathcal{H}}(C_{m},{\mathbf{x}})$ with
${\mathbf{x}}=(1,\zeta_{m}^{k}x,x^{2}:|k|\leq(m-1)/2,\ k\neq\pm m/e)$
by [22, Satz 6.10]. Its specialisation at $x=1$ again has just one non-singleton block (see Section 2.4), containing the linear characters corresponding to the parameters $1,x,x^{2}$. By the description in loc. cit. these correspond to the unipotent principal series characters labelled $1_{W},{\operatorname{sgn}},\boldsymbol{\sigma}_{m/e}$, so Conjecture 4.1 is satisfied.
When $m$ is even, the case when $d=1$ and $e\geq 3$ is entirely similar to the above, so it remains to deal with the case $d=1,e=2$. Here, by [15, Tab. 7.2] the Iwahori–Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,-1))$ specialised at $x=-1$ has just one non-singleton block, containing the four linear characters. By Ennola duality again, the Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,1))$ specialised at $x=1$ also has just one non-singleton block, again consisting of the of the four linear characters. By the description in [22, Def. 6.3] the linear characters in both cases label the same four unipotent characters, confirming Conjecture 4.1 in this case. In fact, this case also follows from the analogue of Lemma 2.7(b). ∎
Note that this result generalises Proposition 3.1, which is for $I_{2}(6)$. There is also a twisted spets ${}^{2}\!I_{2}(m)$, as described in [22, §6]. When $m$ is odd, this can be obtained as an Ennola twist of the untwisted spets of $I_{2}(m)$, so Conjecture 4.1 holds by Proposition 4.3 with Remark 4.2(b). When $m$ is even, this includes the case of the twisted finite reductive groups of types ${}^{2}\!B_{2}$ and ${}^{2}G_{2}$. We have checked that Conjecture 4.1 holds in this case as well.
We next investigate $H_{3}$. In the following we write $e^{\prime},e^{\prime\prime}$ for the two factors of $\Phi_{e}$ over ${\mathbb{Z}}[\sqrt{5}]$, where $e\in\{5,10,15,20,30\}$. For the parameters see [5, Tab. 8.3]. There are four pairs of unipotent characters of $H_{3}$ having the same degree, so only for those is there a choice in the HC-parametrisation.
Proposition 4.4.
Conjecture 4.1 holds for type $H_{3}$ with respect to some complete HC-parametrisation. The block intersections are displayed in Table 5.
Table 5. Block intersections for $H_{3}$
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,3)&\{\phi_{1,0},\phi_{4,3},\phi_{5,2}\},\{\phi_{1,15},\phi_{4,4},\phi_{5,5}\}\\ (1,5^{\prime})&\{\phi_{1,0},\phi_{3,6},\phi_{4,3}\},\{\phi_{1,15},\phi_{3,1},\phi_{4,4}\}\\ (1,10^{\prime})&\{\phi_{1,0},\phi_{1,15},\phi_{3,1},\phi_{3,6}\},\\ &\{H_{2}[\zeta_{5}^{3}]{:}1^{2},H_{2}[\zeta_{5}^{3}]{:}2\},\{H_{2}[\zeta_{5}^{2}]{:}1^{2},H_{2}[\zeta_{5}^{2}]{:}2\}\\ (3,5^{\prime})&\{\phi_{1,0},\phi_{4,3}\},\{\phi_{1,15},\phi_{4,4}\}\\ (3,6)&\{\phi_{1,0},\phi_{5,5}\},\{\phi_{1,15},\phi_{5,2}\}\\ (5^{\prime},5^{\prime\prime})&\{\phi_{1,0},H_{2}[\zeta_{5}^{3}]{:}1^{2}\},\{\phi_{1,15},H_{2}[\zeta_{5}^{2}]{:}2\},\{\phi_{4,3},H_{2}[\zeta_{5}^{2}]{:}1^{2}\},\{\phi_{4,4},H_{2}[\zeta_{5}^{3}]{:}2\}\\ (5^{\prime},10^{\prime})&\{\phi_{1,0},H_{2}[\zeta_{5}^{2}]{:}2\},\{\phi_{1,15},H_{2}[\zeta_{5}^{3}]{:}1^{2}\},\{\phi_{3,1},H_{2}[\zeta_{5}^{2}]{:}1^{2}\},\{\phi_{3,6},H_{2}[\zeta_{5}^{3}]{:}2\}\\ (5^{\prime},10^{\prime\prime})&\{\phi_{1,0},H_{2}[\zeta_{5}^{2}]{:}1^{2}\},\{\phi_{1,15},H_{2}[\zeta_{5}^{3}]{:}2\},\{H_{2}[\zeta_{5}^{3}]{:}1^{2},H_{2}[\zeta_{5}^{2}]{:}2\}\\ \end{array}$
Again, we have omitted all singletons, as well as the pairs
$(1,2),(1,6),(3,10^{\prime})$ where there is just one non-singleton block.
Proof.
The relative Weyl groups and the parameters for the corresponding cyclotomic Hecke algebras are described in [5, Tab. 8.3]. The decomposition matrices for $W(H_{3})$ were computed by Müller [29, §4]. ∎
We now turn to $H_{4}$. There are 29 pairs of characters with equal degree, plus one family of four characters with that property. So for these is there a choice in the HC-parametrisation.
Proposition 4.5.
Conjecture 4.1 holds for type $H_{4}$ with respect to a suitable complete HC-parametrisation.
We refrain again from displaying the block intersections.
Proof.
The relative Weyl groups and the parameters for the corresponding cyclotomic Hecke algebras are described in [5, Tab. 8.3]. The decomposition matrices for $W(H_{3})$ were computed by Müller [29, §4]. The remaining block decompositions were obtained as described in Section 2.6. ∎
4.3. Verifying the spetsial conjecture
Finally, we consider Conjecture 4.1 for some of the primitive spetsial reflection groups $W$.
The degree polynomials of the unipotent characters of the spets associated to the complex reflection group $G_{4}$ are all distinct, so there is a unique complete HC-parametrisation.
Proposition 4.6.
Conjecture 4.1 holds for types $G_{4},G_{6},G_{8}$ and $G_{24}$. The block intersections for $G_{4}$ are displayed in Table 6.
Table 6. Block intersections for $G_{4}$
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{[\phi_{1,0},\phi_{1,4},\phi_{1,8},\phi_{2,3},\phi_{2,1},\phi_{3,2}\},\{Z_{3}[2],Z_{3}[1^{2}]\}\\ (1,3^{\prime})&\{\phi_{1,0},\phi_{1,8}\},\{\phi_{2,5},\phi_{2,1}\}\\ (1,4)&\{\phi_{1,0},\phi_{2,5},\phi_{3,2}\}\\ (1,6^{\prime})&\{\phi_{1,0},\phi_{1,4},\phi_{2,1}\},\{Z_{3}[2],Z_{3}[1^{2}]\}\\ (3^{\prime},3^{\prime\prime})&\{\phi_{1,0},Z_{3}[1^{2}]\},\{\phi_{2,5},Z_{3}[2]\}\\ (3^{\prime},4)&\{\phi_{1,0},\phi_{2,5}\}\\ (3^{\prime},6^{\prime})&\{\phi_{1,0},Z_{3}[2]\},\{\phi_{2,1},Z_{3}[1^{2}]\}\\ (3^{\prime},6^{\prime\prime})&\{\phi_{1,0},Z_{3}[1^{2}]\},\{\phi_{1,8},Z_{3}[2]\}\\ \end{array}$
Again, we have omitted all singletons, as well as pairs that can be obtained
by a Galois automorphism from pairs displayed in the table, like $(1,3^{\prime\prime}),(1,6^{\prime\prime})$.
We also have a partial result for the 5-dimensional group $G_{33}$. Here, again, the dimensions of some irreducible representations are too large to be handled by our methods, and we only obtained approximate block decompositions.
Proposition 4.7.
Conjecture 4.1 holds for type $G_{33}$ except possibly when one of the two Harish-Chandra series is the principal $d$-series, with $d\in\{1,2\}$.
In this case, the 3-dimensional group $G_{26}$ occurs as a relative Weyl group for $d=3,6$.